OpenStudy (anonymous):

What is the value of x? 4(6 + 7) = 4 • 6 + x • 7 x =

alones (alones):

Welcome to OpenStudy ;) 4(6+7)=(4)(6)+x(7) was your problem simplify both sides of the equation. 4(6+7)=(4)(6)+x(7) 52=7x+24 then flip the equation. 7x+24=52 subtract 24 from both sides. 7x+24−24=52−24 7x=28 divide both sides by 7. 7x/7=28/7 x=4

OpenStudy (anonymous):

welcome ill help you

OpenStudy (anonymous):

ok alones did it for me thx